A droplet of ink in an industrial ink-jet printer carries a charge of 1.4×10−10C and is deflected onto paper by a force of 2.4×10−4N.

Find the strength of the electric field (E=F/q)required to produce this force.

Express your answer to two significant figures and include the appropriate units.
Answer #1

Electric field strength, E = F/q = (2.4/1.4)x106 = 1.71 x 106 N/C